Mathematics 241 – Multivariable Calculus Quiz 2 – September 17, 2010

Your Name:

A) (10) Find the area of the parallelogram spanned by u = (3, 2, 1) and v = (1, 1, 2).

B) (10) Find the equation of the plane in R^{3} containing p = (1, 1, 0) and perpendicular to the line {(1, 1, 0) + t(−2, 1, −2) | t ∈ R}.